OpenStudy (hkmiu):

Write an equation for a sine curve that has the given amplitude and period, and which passes through the given point. Amplitude 5, period 6, point (2, 0)

OpenStudy (hkmiu):

y= 5 sin (6x - pi/2) y= 5 sin (3x - 3/2) y= 5 sin ((pi/3)x - pi/2) y= 5 sin ((pi/3)x - (2/3)pi)

OpenStudy (hkmiu):

I dont know how to convert the period and the point

OpenStudy (agent0smith):

\[\large y = a \sin k(x-b)\]they tell you a=5, and the period is 6. \[\large period = \frac{ 2\pi }{ k }\]so find k.

OpenStudy (hkmiu):

is it y= 5 sin (6x - pi/2) ??

OpenStudy (hkmiu):

oh wait no

OpenStudy (hkmiu):

its y= 5 sin ((pi/3)x - (2/3)pi)

OpenStudy (agent0smith):

Once you have k, plug it in the formula, along with a=5, and plug in the point (2,0) \[\large 0 = 5\sin k(2-b)\]find b

OpenStudy (hkmiu):

k is pi/3